How do I graph?
x - 2 = 1/8 (y + 1)^2

Answer by Alan3354(69443)


x - 2 = 1/8 (y + 1)^2
x = 2 + 1/8 (y + 1)^2
Pick values for y and find x
Plot the points
Draw a curve thru them.